(i) C(X +Y,Z)= C(X,Z)+C(Y,Z), where X and Y are two random vectors of the same dimension and C(X,Z)=C(Z,X).

(ii) If m=n, V(X+Y) =V(X) + V(Y)+ 2C(X,Y).


1
Expert's answer
2021-09-27T15:30:30-0400

Let

"X=\left(X_1 ,....X_n \right)"

"Y=\left(Y_1 ,....Y_n \right)"

"Z=\left(Z_1 ,....Z_n \right)"

be any random vectors on the same probability space "P=(\varOmega, \varSigma, \mu)"

Then covatiance matrix C(X+Y,Z)=

"E\left( (X+Y-E(X+Y))(Z-EZ)^T \right)=\\
E\left( (X+Y-EX-EY))(Z-EZ) ^T \right)=\\
E\left( (X-EX)+(Y-EY))(Z-EZ)^T \right)=\\
E\left( (X-EX)(Z-EZ)^T + (Y-EY)(Z-EZ)^T \right)=\\
E\left( (X-EX)(Z-EZ)^T \right)+E\left( (Y-EY)(Z-EZ)^T \right)=\\
C(X,Z)+C(Y,Z)"

​First identity is proved.

During proving property of random matrices V,W : E(V+W)=EV+EW have being applied.

2) For proving second statement we need in additional condition C(X,Y)=C(Y,X)

"V(X+Y)=C(X+Y,X+Y)=]by\space 1]=C(X,X)+C(X,Y)+C(Y,X)+C(Y,Y)=\\
=V(X)+V(Y)+C(X,Y)+C(X,Y)^T=\\
[by\space asymption \space C(X,Y)=C(Y,X)]= \\
V(X)+V(Y)+2\cdot C(X,Y)"